# Models, Methods and Waveforms for Estimation and Prediction of Doubly Sparse Time-Varying Channels

### v1

@@ V Numerical results
We first present the results pertaining to the estimation of DS-LTV channels with the on-grid approximation of Definition 1. AFDM sparse recovery performance is compared to that of OFDM and OTFS. We used 100 realizations of channels having a Type-1 delay-Doppler sparsity with $p_{\rm d}=0.2$, $p_{\rm D}\in\{0.2,0.4\}$ and $N=4096,L=30,Q=7$ (corresponding to a 30 MHz transmission at a 70 GHz carrier frequency, a maximum target moving speed of 396 km/h and a maximum target range of 300 meters). For both OFDM and AFDM, sparse recovery of $\boldsymbol{\alpha}$ is done using HiHTP (Algorithm 1) based on the pilot pattern described in Figures 3 and 4, respectively. For OTFS, since sensing is done without compression, non-compressive estimation algorithms can be used [15]. For each waveform, the pilot overhead was set in such a way that the mean squared error ${\rm MSE}\triangleq\mathbb{E}[\left\|\hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}\right\|^{2}]$ is approximately $10^{-4}$ at $\mathrm{SNR}=20$ dB. Fig. 8 shows an advantage of AFDM in terms of pilot overhead i.e., the number of samples in each frame needed as pilots and guards to achieve the target MSE performance. This performance validates the theoretical result given in Theorems 1 and 2 implying that the overhead needed to get the same HiRIP level is smaller for AFDM compared to SCM and OFDM.
Refer to caption Refer to caption
(a) $p_{\rm D}=0.2$ (b) $p_{\rm D}=0.4$
Figure 8: MSE and pilot overhead for $N=4096,L=30,Q=7,p_{\rm d}=0.2$, $N_{\rm ofdm,symb}=16$, $N_{\rm otfs}=16$, $M_{\rm otfs}=256$. Overhead: $N_{\rm p,td}N_{\rm p,fd}+(N_{\rm ofdm,symb}-1)(L-1)$ for OFDM, $\min(4Q+1,N_{\rm otfs})\min(2L-1,M_{\rm otfs})$ for OTFS, $N_{\rm p}\left((L-1)P+1\right)+(L-1)P+4Q$ for AFDM.
@@
Figure 9: MSE as function of $Q_{\rm BEM}$ for the estimation of a channel tap with one cluster of fractional Doppler shifts $(N=2048,Q=7,N_{\rm p}=46)$ for two SNR values (a) $\text{SNR}=10$ dB, (b) $\text{SNR}=40$ dB
We now turn to the estimation and extrapolation of DS-LTV channels with off-grid Doppler shifts as modeled by Definition 4. We start with the configuration of our BEM model proposed in Subsection IV-C. The effect of the order $Q_{\rm BEM}$ of the elementary BEM on the MSE with different extrapolation horizons, namely $N_{\rm ext}=1000$, $N_{\rm ext}=500$ and $N_{\rm ext}=0$ (marked as CE), is shown in Figure 9. While increasing $Q_{\rm BEM}$ improves the precision of the model, it increases the number of unknown to be estimated revealing the trade-off between model precision and estimation performance. This is why $Q_{\rm BEM}=3$ yields the best MSE at the lower SNR value as opposed to $Q_{\rm BEM}=4$ at the higher. Note that both of these two $Q_{\rm BEM}$ values being small confirm the result given in Theorem 3.
Figure 10 compares the MSE performance of our method based on multiple shifted elementary BEMs with $Q_{\rm BEM}=4$ to two other approaches, namely the single-BEM approach of (39) (configured such that $\frac{\check{Q}_{\rm BEM}}{\textrm{\# active Doppler shifts per delay tap}}=4$), and a refined-grid approach [6] (with a refinement factor of $O=4$). The comparison is done for both channel estimation and prediction assuming $N_{\rm ext}=500$ and the underlying waveform is AFDM with $P=1$. The values of $Q_{\rm BEM}$, $\check{Q}_{\rm BEM}$, and $O$ are chosen equal to ensure a fair comparison across all methods, both in terms of model size (number of unknown parameters) and computational complexity.
Figure 10: MSE performance of AFDM using different approaches ($N=2048,L=20,Q=7,p_{\rm d}=0.2,p_{\rm D}=0.2$)
For channel estimation, the multiple shifted BEMs approach and the single-BEM method achieve similar performance, both outperforming the refined-grid technique (due to the latter being ill-conditioned since $N_{\rm D}>1$ as explained at the beginning of Section IV). However, for channel prediction, our approach exhibits a slight performance degradation compared to the single BEM method. Nevertheless, as highlighted in Remark 3, our method requires a significantly smaller codebook size than the single BEM approach shown by Table I (for the same setting as in Figures 11 and 10), making it more efficient in practical implementations.
TABLE I: Size of codebook for each method
Method Codebook Size
One BEM 2,621,360
Multiple BEMs 1,200
Finally, we shift our attention to comparing the sparse recovery performance of AFDM with that of OFDM. We used 100 realizations of channels having a off-grid Type-1 delay-Doppler sparsity with $p_{\rm d}=0.2$, $p_{\rm D}=0.2$ and $N=2048,L=20,Q=7$ (corresponding to a 15 MHz transmission at a 70 GHz carrier frequency, a maximum target moving speed of 396 km/h and a maximum target range of 400 meters). As highlighted previously in remark 2, Figure 11 illustrates that, when dealing with off-grid Doppler shift values, AFDM maintains its superiority proved with the on-grid model. Indeed, the values of $N_{\rm p}$, $N_{\rm p,t}$ and $N_{\rm p,f}$ used to generate the figure, while yielding almost identical pilot overheads for OFDM and AFDM ($\textrm{overhead}_{\rm AFDM}=767$, $\textrm{overhead}_{\rm OFDM}=769$, result in a superior MSE performance for AFDM in both channel estimation and channel prediction tasks.
Figure 11: MSE of Channel estimation and prediction for $N=2048,L=20,Q=7,p_{\rm d}=0.2$, $p_{\rm D}=0.2$,$N_{\rm s}=32$, $N_{\rm fft}$=64
VI Conclusions
This paper investigated channel estimation for doubly sparse LTV channels, addressing both on-grid and off-grid Doppler shifts. The on-grid approximation was used to provide rigorous recovery guarantees when channel estimation is done based on different underlying waveforms. To handle off-grid effects, we introduced a DPSS-based model consisting of multiple elementary BEMs and we provided analytical precision guarantees for that model. Additionally, a channel prediction method was proposed by taking advantage of DPSS extrapolation capability. Numerical results confirmed AFDM’s superiority in both estimation and prediction under double (delay-Doppler) sparsity, making it a strong candidate for future wireless systems.
Appendix A Proof of Theorem 1
In the case of SCW, defining $\tilde{\alpha}_{l,q}\triangleq\alpha_{l,q}e^{\imath 2\pi\frac{lq}{N}}$, rearranging $\mathbf{y}_{\rm p}$ into $\tilde{\mathbf{y}}_{\rm p}$ (composed of $L$ successive blocks, with the $l$-th blocks composed of the $l$-th sample in each of the $N_{\rm p}$ pilot intervals) and assuming $12Q+$ 1 divides $N$ and that the $p$-th pilot position (for any $p\in\llbracket 1\mkern 1.5mu..\mkern 1.5muN_{\rm p}\rrbracket$) satisfies $m_{p}=q_{p}\frac{N}{2Q+1}$ for some $q_{p}\in\llbracket 0\mkern 1.5mu..\mkern 1.5mu2Q\rrbracket$, we can write $\tilde{\mathbf{y}}_{\rm p}=\widetilde{\mathbf{M}}_{\rm p}^{\rm scw}\tilde{\boldsymbol{\alpha}}$ where $\tilde{\boldsymbol{\alpha}}$ is the vectorized form of $\tilde{\alpha}_{l,q}$ and $\widetilde{\mathbf{M}}_{\rm p}^{\rm scw}\triangleq\mathbf{I}_{L}\otimes\left(\mathrm{diag}\left(\mathrm{p}_{1},\ldots,\mathrm{p}_{N_{\rm p}}\right)\overline{\mathbf{F}}_{2Q+1,N_{\rm p}}\right)$. Here, $\overline{\mathbf{F}}_{2Q+1,N_{\rm p}}$ is the partial inverse Fourier measurement matrix formed from $N_{\rm p}$ rows of the $(2Q+1)$-point inverse DFT matrix. The HiRIP of $\widetilde{\mathbf{M}}_{\rm p}^{\rm scw}$ can thus be derived and proven to be equal to the value given in the theorem statement by using the known RIP of partial inverse Fourier measurement matrices [28, Theorem 4.5] followed by applying [18, Theorem 4] pertaining to the HiRIP of hierarchical measurement matrices having the Kronecker property. This completes the part of the proof related to SCM.
As for OFDM, it can be shown that the estimation problem has a measurement matrix $\widetilde{\mathbf{M}}_{\rm p}^{\rm ofdm}\triangleq\left(\mathrm{diag}\left(\mathrm{p}_{1},\ldots,\mathrm{p}_{N_{\rm p,f}}\right)\mathbf{F}_{L,N_{\rm p,f}}\right)\otimes\overline{\mathbf{F}}_{2Q+1,N_{\rm p,t}}$ where $\mathbf{F}_{L,N_{\rm p,f}}$ is the partial Fourier measurement matrix formed from $N_{\rm p,f}$ rows of the $L$-point inverse DFT matrix. The value of the HiRIP of $\widetilde{\mathbf{M}}_{\rm p}^{\rm ofdm}$ given in the statement of the theorem thus follows from the RIP of the partial Fourier measurement matrix and the HiRIP result pertaining to Kronecker hierarchical measurements.
Appendix B Proof of Theorem 2
First, out of the pilot samples set $\mathcal{P}$, consider the subset $\mathcal{P}_{p}$ associated with the $p$-th pilot symbol transmitted at the DAFT index $m_{p}$ (Fig. 4). To homogenize the sensing signal model associated with edge samples and inner samples of $\mathcal{P}_{p}$, we apply two overlap-add operations: adding the samples received within the index interval $\llbracket m_{p}-Q\mkern 1.5mu..\mkern 1.5mum_{p}-1\rrbracket$ to those received within $\llbracket m_{p}+(L-1)P-Q\mkern 1.5mu..\mkern 1.5mum_{p}+(L-1)P-1\rrbracket$ and the samples received within $\llbracket m_{p}+(L-1)P+1\mkern 1.5mu..\mkern 1.5mum_{p}+(L-1)P+Q\rrbracket$ to those received within $\llbracket m_{p}+1\mkern 1.5mu..\mkern 1.5mum_{p}+Q\rrbracket$. Now, define
$\mathcal{D}_{l}\triangleq\left\{(\tilde{l},q)\mathrm{s.t.}(q+P\tilde{l})_{(L-1)P+1}=l\right\}$ (43)
as the set of delay-Doppler grid points that potentially contribute to the pilot sample received at DAFT domain index $l\in\llbracket m_{p}\mkern 1.5mu..\mkern 1.5mum_{p}+(L-1)P\rrbracket$ (Fig. 12) after the two overlap-add operations described above. Note that $\mathcal{D}_{l}$ does not depend on the pilot symbol index $p$ and that it has a cardinality that does not change with $l$ and which satisfies $\left|\mathcal{D}_{l}\right|\leq 2\lceil\frac{Q}{P}\rceil+1$. Next, define $\boldsymbol{\alpha}_{\mathcal{D}_{l}}\triangleq\left[\alpha_{l,q}\right]_{(l,q)\in\mathcal{D}_{l}}$ and $\widetilde{\boldsymbol{\alpha}}\triangleq\left[\boldsymbol{\alpha}_{\mathcal{D}_{0}}^{\rm T}\quad\cdots\quad\boldsymbol{\alpha}_{\mathcal{D}_{(L-1)P}}^{\rm T}\right]^{\rm T}$. The entries of $\widetilde{\boldsymbol{\alpha}}$ are just a permutation of the entries of $\boldsymbol{\alpha}$ and estimating one of these vectors directly gives an estimate of the other. Now, it can be shown that when we set $P$ as in the theorem and $\epsilon>0$ as small as needed, then $\widetilde{\boldsymbol{\alpha}}$ is $\left(\tilde{s}_{\rm d},\tilde{s}_{\rm D}\right)$-hierarchically sparse with high probability.
Refer to caption Refer to caption
(a) (b)
@@ Appendix B Proof of Theorem 2
Figure 12: Two examples of the set $\mathcal{D}_{l}$ ($P=1$, $m_{p}=0$) (a) for an $l$ resulting in a whole diagonal, (b) for an $l$ resulting in a wrapped diagonal. In each one of the two examples, the grid points forming $\mathcal{D}_{l}$ are shown surrounded by red rings.
$\tilde{s}_{\rm d}=(L-1)P+1,\quad\tilde{s}_{\rm D}=(1+\epsilon)\log(LP)\>.$ (44)
@@
Indeed, the first level (of size $(L-1)P+1$) of $\widetilde{\boldsymbol{\alpha}}$ is sensed without compression with a number of measurements equal to $(L-1)P+1$ while $\tilde{s}_{\rm D}$ can be determined thanks to Definition 1 and Assumptions 1 and 2 and applying the same approach as in the proof of Lemma 1 to $\tilde{S}_{{\rm D},l}\triangleq\sum_{(\tilde{l},q)\in\mathcal{D}_{l}}I_{\tilde{l},q}$. Now, we can write the signal model of sensing $\widetilde{\boldsymbol{\alpha}}$ as

### v2

@@ V Numerical results
We first present the results pertaining to the estimation of DS-LTV channels with the on-grid approximation of Definition 1. AFDM sparse recovery performance is compared to that of OFDM and OTFS. Note that OFDM pilot overhead results can be indicative of the pilot overhead performance of SCM and other waveforms using time domain embedded pilots such as IFDM. We used 100 realizations of channels having a Type-1 delay-Doppler sparsity with $p_{\rm d}=0.2$, $p_{\rm D}\in\{0.2,0.4\}$ and $N=4096,L=30,Q=7$ (corresponding to a 30 MHz transmission at a 70 GHz carrier frequency, a maximum target moving speed of 396 km/h and a maximum target range of 300 meters, or equivalently to 10 MHz transmission at a 30 GHz carrier frequency, a maximum speed of 57 km/h and a maximum target range of 900 meters). For both OFDM and AFDM, sparse recovery of $\boldsymbol{\alpha}$ is done using HiHTP (Algorithm 1) based on the pilot pattern described in Figures 3 and 4, respectively. For OTFS, since sensing is done without compression, non-compressive estimation algorithms can be used based on the pilot pattern of Fig. 5 [20]. For each waveform, the pilot overhead was set in such a way that the mean squared error ${\rm MSE}\triangleq\mathbb{E}[\left\|\hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}\right\|^{2}]$ is approximately $10^{-4}$ at $\mathrm{SNR}=20$ dB. Fig. 9 shows an advantage of AFDM in terms of pilot overhead, i.e., the number of samples in each frame needed as pilots and guards to achieve the target MSE performance. This performance validates the theoretical result, given in Theorems 1 and 2, implying that the overhead needed to achieve the same HiRIP level is smaller for AFDM compared to SCM and OFDM.
(a) $p_{\rm D}=0.2$
(b) $p_{\rm D}=0.4$
Figure 9: On-grid sparse-recovery MSE and pilot overhead for $N=4096,L=30,Q=7,p_{\rm d}=0.2$, $N_{\rm ofdm,symb}=16$, $N_{\rm otfs}=16$, $M_{\rm otfs}=256$ (overhead comparison done at comparable MSE levels for the three waveforms using (24), (25) and (26))
In Fig. 10, we compare the pilot overhead (computed as in Fig. 9) needed by the three waveforms in four different configurations of the delay and Doppler sparsity level parameters, namely $(s_{\rm d},s_{\rm D})=(0.1,0.2)$, $(s_{\rm d},s_{\rm D})=(0.2,0.2)$, $(s_{\rm d},s_{\rm D})=(0.2,0.4)$ and $(s_{\rm d},s_{\rm D})=(0.4,0.2)$. Note that AFDM, configured with $P_{\rm afdm}=1$ for the first scenario and with $P_{\rm afdm}=2$ for the remaining ones, has the smallest overhead in all four scenarios. Also note that for OTFS, the overhead is the same in the four scenarios, as its pilot resources are determined as a function of the maximum delay and Doppler spreads, and not as a function of sparsity levels. This also hints at the possibility that the performance advantage over OTFS tends to diminish as the channel becomes less sparse. This intuition is validated by the figure, which shows a smaller gap in overhead values in the fourth sparsity scenario. The bit error rate (BER) and spectral efficiency (SE) of the considered waveforms in the first and fourth sparsity scenarios are given in Fig. 11 assuming quadrature quadrature phase-shift keying (QPSK) data symbols. While OTFS achieves the best BER, as it is a full-diversity waveform, AFDM can achieve a very close BER performance by adjusting its chirp parameter $c_{1}=\frac{P_{\rm afdm}}{2N}$ to the sparsity level. However, the gap in performance is much wider in favor of AFDM when pilot overhead is considered while evaluating spectral efficiency.
Figure 10: Pilot overhead for sparse recovery ($N=2048,L=30,Q=7$)
(a) BER
(b) SE
Figure 11: BER and SE comparison ($N=2048,L=30,Q=7,p_{\rm D}=0.2$)
(a) $\text{SNR}=10$ dB
(b) $\text{SNR}=40$ dB
Figure 12: Effect of parameter $Q_{\rm BEM}$ of the “multiple shifted elementary BEMs” method on the MSE performance of both off-grid LTV channel estimation and prediction ($N=2048,Q=7,N_{\rm p}=46$)
We now turn to the estimation and extrapolation of DS-LTV channels with off-grid Doppler shifts, as modeled by Definition 4. We start with the configuration of the parameters of our BEM model proposed in Subsection IV-C. The effect of the order $Q_{\rm BEM}$ of the elementary BEM on the MSE with different extrapolation horizons, namely $N_{\rm ext}=1000$, $N_{\rm ext}=500$ and $N_{\rm ext}=0$ (marked as “channel estimation”), is shown in Figure 12. While increasing $Q_{\rm BEM}$ improves the precision of the model, it also increases the number of unknowns to be estimated, revealing the trade-off between model precision and pilot overhead. This is why $Q_{\rm BEM}=3$ yields the best MSE at the lower SNR value as opposed to $Q_{\rm BEM}=4$ at the higher. Note that the fact that both $Q_{\rm BEM}$ values are small confirms the result given in Theorem 3.
Figure 13 compares the MSE performance of the “multiple shifted elementary BEMs” method, with $Q_{\rm BEM}=4$, to two other approaches, namely the multi-band single-BEM approach of (49) (configured such that $\frac{\check{Q}_{\rm BEM}}{\textrm{\# active Doppler shifts per delay tap}}=4$), and a refined-grid approach [9] (with a refinement factor of $O=4$). Note that these values of $Q_{\rm BEM}$, $\check{Q}_{\rm BEM}$, and $O$ are chosen to be equal to ensure a fair comparison across all methods, both in terms of model size (number of unknown parameters) and computational complexity. The comparison is conducted for both channel estimation (CE) and prediction, assuming $N_{\rm ext}=500$ and that the underlying waveform is AFDM with $P_{\rm afdm}=1$.
Figure 13: MSE of different off-grid LTV channel estimation and prediction methods under AFDM ($N=2048,L=20,Q=7,p_{\rm d}=p_{\rm D}=0.2$)
For channel estimation, the multiple shifted BEMs approach and the multi-band single-BEM method achieve similar performance, both outperforming the refined-grid technique (due to the latter being ill-conditioned since $N_{\rm D}>1$ as explained at the beginning of Section IV). However, for channel prediction, our approach exhibits a slight performance degradation compared to the single BEM method. Nevertheless, as highlighted in Remark 7, our method requires a significantly smaller codebook size than the multi-band single-BEM approach shown by Table I (for the same setting as in Figures 15 and 13), making it more efficient in practical implementations.
TABLE I: Size of codebook for each method ($L=20,Q=7,p_{\rm d}=0.2,p_{\rm D}=0.2$)
Method Codebook Size
Multi-band single BEM [31]. Number of Doppler bands = $(2Q+1)p_{\rm D}=3$, $\check{Q}_{\rm BEM}=12$ 2,621,360
Multiple shifted BEMs (36). Number of elementary BEMs = $(2Q+1)p_{\rm D}=3$, $Q_{\rm BEM}=4$ 1,200
To test the effect of mobility on the proposed channel estimation and prediction methods when using AFDM, we use two maximum Doppler shift values: $Q=3$ and $Q=7$, corresponding to maximum speeds of 170 km/h and 396 km/h, respectively. Doppler sparsity is $p_{\rm D}=0.4$ in the lower-mobility case and $p_{\rm D}=0.2$ in the other case, ensuring that the number of active Doppler clusters is the same in both scenarios. As expected, the MSE performance is better in the lower-mobility case. However, the degradation in performance is limited when the maximum speed is almost doubled, demonstrating the robustness of both AFDM and the proposed estimation and prediction schemes under high mobility.
Figure 14: Effect of mobility (value of $Q$) on MSE performance of off-grid LTV channel estimation and prediction using the “multiple shifted elementary BEMs” and DPSS extrapolation approaches under AFDM ($N=2048,L=20,Q_{\rm BEM}=4,p_{\rm d}=0.2$)
Finally, we demonstrate that the superiority of AFDM in terms of sparse-recovery performance, as established using compressed-sensing tools under the on-grid model, extends to the off-grid model and LMMSE channel estimation. We used 100 realizations of channels having a off-grid Type-1 delay-Doppler sparsity with $p_{\rm d}=0.2$, $p_{\rm D}=0.2$ and $N=2048,N_{\rm ofdm,symb}=32,L=20,Q=7$. As highlighted in Remark 5, Fig. 15 illustrates that when dealing with off-grid Doppler shifts, AFDM maintains the superiority that was proven with the on-grid model. Indeed, the values of $N_{\rm p}$, $N_{\rm ofdm,symb}$, $N_{\rm p,t}$ and $N_{\rm p,f}$ used to generate the figure, while yielding almost identical pilot overheads for OFDM and AFDM ($\textrm{overhead}_{\rm AFDM}=766$, $\textrm{overhead}_{\rm OFDM}=788$, result in a superior MSE performance for AFDM in both channel estimation and channel prediction tasks.
Figure 15: MSE of off-grid LTV channel estimation and prediction using AFDM and OFDM for $N=2048,L=20,Q=7,p_{\rm d}=p_{\rm D}=0.2$$N_{\rm ofdm,symb}=32$, $N_{\rm fft}=64$ (MSE comparison done at comparable overhead levels for the two waveforms)
VI Conclusions
This paper investigated channel estimation for doubly sparse LTV channels, addressing both on-grid and off-grid Doppler shifts. The on-grid approximation, through its link to the hierarchical-sparsity compressed-sensing paradigm, was used to provide a rigorous performance comparison when channel estimation is conducted based on different underlying waveforms. This theoretical comparison established that the AFDM waveform can achieve the same sparse-recovery guarantees while requiring reduced pilot overhead compared to SCM, OFDM, and OTFS. To relax the on-grid assumption, we introduced a DPSS-based channel representation model consisting of multiple frequency-shifted elementary BEMs capable of handling off-grid Doppler effects, and we complemented it with analytical precision guarantees. We next used the proposed channel representation to conceive a practical LMMSE estimator for LTV channels. Additionally, a channel predictor was introduced by taking advantage of the DPSS extrapolation capability of the proposed LMMSE estimator. Numerical results confirmed AFDM superiority in both estimation and prediction under delay-Doppler sparsity, making it a strong candidate for future wireless systems.
Appendix A Proof of Theorem 1
In the case of SCW, defining $\tilde{\alpha}_{l,q}\triangleq\alpha_{l,q}e^{\imath 2\pi\frac{lq}{N}}$, rearranging $\mathbf{y}_{\rm p}$ into $\tilde{\mathbf{y}}_{\rm p}$ (composed of $L$ successive blocks, with the $l$-th blocks composed of the $l$-th sample in each of the $N_{\rm p}$ pilot intervals) and assuming $2Q+1$ divides $N$ and that the $p$-th pilot position (for any $p\in\llbracket 1\mkern 1.5mu..\mkern 1.5muN_{\rm p}\rrbracket$) satisfies $m_{p}=q_{p}\frac{N}{2Q+1}$ for some $q_{p}\in\llbracket 0\mkern 1.5mu..\mkern 1.5mu2Q\rrbracket$, we can write $\tilde{\mathbf{y}}_{\rm p}=\widetilde{\mathbf{M}}_{\rm p}^{\rm scw}\tilde{\boldsymbol{\alpha}}$ where $\tilde{\boldsymbol{\alpha}}$ is the vectorized form of $\tilde{\alpha}_{l,q}$ and
$\widetilde{\mathbf{M}}_{\rm p}^{\rm scw}\triangleq\mathbf{I}_{L}\otimes\left(\mathrm{diag}\left(\mathrm{p}_{1},\ldots,\mathrm{p}_{N_{\rm p}}\right)\overline{\mathbf{F}}_{2Q+1,N_{\rm p}}\right)$ (53)
Here, $\overline{\mathbf{F}}_{2Q+1,N_{\rm p}}$ is the partial inverse Fourier measurement matrix formed from $N_{\rm p}$ rows of the $(2Q+1)$-point inverse DFT matrix. The HiRIP of $\widetilde{\mathbf{M}}_{\rm p}^{\rm scw}$ can thus be derived and proven to be equal to the value given in the theorem statement by using the known RIP of partial inverse Fourier measurement matrices [37, Theorem 4.5] followed by applying [24, Theorem 4] pertaining to the HiRIP of hierarchical measurement matrices having the Kronecker property. This completes the part of the proof related to SCM. As for OFDM, the measurement matrix of the estimation problem is
$\widetilde{\mathbf{M}}_{\rm p}^{\rm ofdm}\triangleq(\mathrm{diag}(\mathrm{p}_{1},\ldots,\mathrm{p}_{N_{\rm p,f}})\mathbf{F}_{L,N_{\rm p,f}})\otimes\overline{\mathbf{F}}_{2Q+1,N_{\rm p,t}}$ (54)
where $\mathbf{F}_{L,N_{\rm p,f}}$ is the partial Fourier measurement matrix formed from $N_{\rm p,f}$ rows of the $L$-point inverse DFT matrix. The HiRIP of $\widetilde{\mathbf{M}}_{\rm p}^{\rm ofdm}$ thus follows from the RIP of the partial Fourier measurement matrix and the HiRIP result pertaining to Kronecker hierarchical measurements.
Appendix B Proof of Theorem 2
First, out of the pilot samples set $\mathcal{P}$, consider the subset $\mathcal{P}_{p}$ associated with the $p$-th pilot symbol transmitted at the DAFT index $m_{p}$ (Fig. 4). To homogenize the sensing signal model associated with edge samples and inner samples of $\mathcal{P}_{p}$, we apply two overlap-add operations: adding the samples received within the index interval $\llbracket m_{p}-Q\mkern 1.5mu..\mkern 1.5mum_{p}-1\rrbracket$ to those received within $\llbracket m_{p}+(L-1)P_{\rm afdm}-Q\mkern 1.5mu..\mkern 1.5mum_{p}+(L-1)P_{\rm afdm}-1\rrbracket$ and the samples received within $\llbracket m_{p}+(L-1)P_{\rm afdm}+1\mkern 1.5mu..\mkern 1.5mum_{p}+(L-1)P_{\rm afdm}+Q\rrbracket$ to those received within $\llbracket m_{p}+1\mkern 1.5mu..\mkern 1.5mum_{p}+Q\rrbracket$. Now, define
$\mathcal{D}_{l}\triangleq\left\{(\tilde{l},q)\mathrm{s.t.}(q+P_{\rm afdm}\tilde{l})_{(L-1)P_{\rm afdm}+1}=l\right\}$ (55)
as the set of delay-Doppler grid points that potentially contribute to the pilot sample received at DAFT domain index $l\in\llbracket m_{p}\mkern 1.5mu..\mkern 1.5mum_{p}+(L-1)P_{\rm afdm}\rrbracket$ (Fig. 16) after the two overlap-add operations described above. Note that $\mathcal{D}_{l}$ does not depend on the pilot symbol index $p$ and that it has a cardinality that does not change with $l$ and which satisfies $\left|\mathcal{D}_{l}\right|\leq 2\lceil\frac{Q}{P_{\rm afdm}}\rceil+1$. Next, define $\boldsymbol{\alpha}_{\mathcal{D}_{l}}\triangleq\left[\alpha_{l,q}\right]_{(l,q)\in\mathcal{D}_{l}}$ and $\widetilde{\boldsymbol{\alpha}}\triangleq[\boldsymbol{\alpha}_{\mathcal{D}_{0}}^{\rm T}\quad\cdots\quad\boldsymbol{\alpha}_{\mathcal{D}_{(L-1)P_{\rm afdm}}}^{\rm T}]^{\rm T}$. The entries of $\widetilde{\boldsymbol{\alpha}}$ are just a permutation of $\boldsymbol{\alpha}$ and estimating one of these vectors directly provides an estimate of the other. Now, it can be shown that when we set $P_{\rm afdm}$ as in the theorem and $\epsilon>0$ as small as needed, then $\widetilde{\boldsymbol{\alpha}}$ is $\left(\tilde{s}_{\rm d},\tilde{s}_{\rm D}\right)$-hierarchically sparse with high probability:
@@ Appendix B Proof of Theorem 2
Figure 16: Two examples of the set $\mathcal{D}_{l}$ ($P_{\rm afdm}=1$, $m_{p}=0$) (a) for an $l$ resulting in a whole diagonal, (b) for an $l$ resulting in a wrapped diagonal. The grid points forming $\mathcal{D}_{l}$ are shown surrounded by red rings.
$\tilde{s}_{\rm d}=(L-1)P_{\rm afdm}+1,\quad\tilde{s}_{\rm D}=(1+\epsilon)\log(LP_{\rm afdm})\>.$ (56)
@@
Indeed, the first level (of size $(L-1)P_{\rm afdm}+1$) of $\widetilde{\boldsymbol{\alpha}}$ is sensed without compression, with a number of measurements equal to $(L-1)P_{\rm afdm}+1$ while $\tilde{s}_{\rm D}$ can be determined thanks to Definition 1 and Assumptions 1 and 2 and applying the same approach as in the proof of Lemma 1 to $\tilde{S}_{{\rm D},l}\triangleq\sum_{(\tilde{l},q)\in\mathcal{D}_{l}}I_{\tilde{l},q}$. Now, we can write the signal model of sensing $\widetilde{\boldsymbol{\alpha}}$ as